OpenStudy (anonymous):

How many real-number solutions does the equation have? 5. –7x² + 6x + 3 = 0 one solution two solutions no solutions infinitely many solutions

OpenStudy (solomonzelman):

plug the numbers into the quadratic formula and solve can you do that?

OpenStudy (anonymous):

I forget how

OpenStudy (solomonzelman):

\[\huge\color{blue}{ -7 }x^2+\huge\color{blue}{ 6 }x+\huge\color{blue}{ 3 }=0\]

OpenStudy (anonymous):

thnx

OpenStudy (solomonzelman):

See your a b c ? now, \[\huge\color{red}{ \frac{-b±\sqrt{b^2-(4 \times a \times c)}}{2a} }\]